OpenStudy (anonymous):

Select the best possible first step to solving the system by first eliminating the y variable. 4x – 5y = 4 3x + 2y = –20 A. None of the above B. Multiply the first equation by 2 and multiply the second equation by 5. C. Multiply the first equation by 2 and multiply the second equation by –5. D. Multiply the first equation by –2 and multiply the second equation by 5.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

\[4x-5y=4 \rightarrow \times2\]

OpenStudy (anonymous):

\[3x+2y=-20\rightarrow \times5\]

OpenStudy (anonymous):

B. Multiply the first equation by 2 and multiply the second equation by 5. 4x – 5y = 4 3x + 2y = –20 -------------- 2*(4x – 5y = 4 ) 5*(3x + 2y = –20) ---------------- 8x – 10y = 8 15x + 10y = –100

OpenStudy (anonymous):

eliminate the y \[-10y +10y\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

8x – 10y = 8 15x + 10y = –100 ---------------------- 23x=-92

OpenStudy (anonymous):

so the y can be eliminated

OpenStudy (anonymous):

thus y is eliminated

OpenStudy (anonymous):

and x=-92 div 23

jhonyy9 (jhonyy9):

b is right